# On the Convergence of Muon and Beyond

### v2

@@ 4 Convergence Analysis
We begin our analysis by situating it within the established context of first-order stochastic optimization. The $\mathcal{O}(T^{-1/4})$ convergence rate is a well-known bottleneck for methods like SGD, and recent work by Shen et al. (2025) confirmed this limitation for the standard Muon algorithm. Accordingly, our first step in this section is to analyze the Muon-MVR1 variant and formally reaffirm this baseline convergence rate in Theorem 4.1. While the rate itself is standard, our primary contribution here lies in the novel and unified analytical framework we develop to prove it, as this framework will be instrumental for subsequent results.
@@
These assumptions are quite common Zhou et al. (2018); Chen et al. (2018); Huang et al. (2021); Guo et al. (2021); Li et al. (2023); Wang et al. (2023); Xie et al. (2024); Yuan et al. (2024).
@@ 4.1.1 Option MVR1
Suppose Assumptions 4.1, 4.2, and 4.4 hold. Consider Algorithm 1 with a learning rate of $\eta_{t}=t^{-3/4}$. The expected average squared norm of the gradient is bounded for the following options:
@@
Theorems 4.1 establishes that as $T\to\infty$, the leading terms diminish to zero, and the algorithm converges to a neighborhood of a stationary point. The size of this neighborhood is determined by the stochastic gradient variance, the learning rate, and the momentum parameter. Notably, this convergence bound is free of any non-vanishing additive error term that depends on the dimension $n$ Sato et al. (2025). We absorb the dimensional dependence into the $\mathcal{O}$-notation to define $\tilde{\mathcal{O}}$, which yields the following complexity for the algorithm:
$\min_{t=1,\dots,T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq\tilde{\mathcal{O}}(T^{-1/4})$
@@ 4.1.2 Option MVR2
While prior work has established a convergence rate of $\tilde{\mathcal{O}}(T^{-1/3})$ for variance-reduction in non-convex settings for various algorithms, such as SGD Fang et al. (2018); Cutkosky & Orabona (2019); Zhou et al. (2020) and Adam Huang et al. (2021); Yuan et al. (2024), the theoretical convergence properties of this technique when applied to the Muon optimizer have remained an open question. We bridge this theoretical gap in Theorem 4.2, which rigorously proves that Option MVR2 of Algorithm 1 (Muon-MVR2) achieves the same $\tilde{\mathcal{O}}(T^{-1/3})$ convergence rate in the general non-convex setting.
@@
Under Assumptions 4.1, 4.3, and 4.4, consider Algorithm 1 with Option MVR2 and update rule (5) parameters set as $\beta_{t}=1-\eta_{t}$, $\eta_{t}=t^{-2/3}$ for $t\geq 1$, and $\gamma=1$. Then, the following bounds hold:
[…]
@@
Theorem 4.2 establishes a significant theoretical advantage. In contrast to prior work Sfyraki & Wang (2025) that requires a computationally expensive growing batch size ($b=\mathcal{O}(T^{1/3})$) to control gradient variance, our result shows that Muon-MVR2 achieves the same $\tilde{\mathcal{O}}(T^{-1/3})$ convergence rate in a more practical, fixed mini-batch setting. This is made possible by the specific hyperparameter schedule, where the learning rate $\eta_{t}=t^{-2/3}$ and the momentum parameter $\beta_{t}=1-\eta_{t}$ are critically coupled. This schedule ensures a careful balance between optimization progress and noise suppression.
@@
As noted in Yuan et al. (2024), a more sophisticated, adaptive setting for the variance-reduction parameter $\gamma$ can be employed. Specifically, by setting $\gamma=\gamma_{t}=1-\frac{A_{t}}{\beta_{t}}$, where $A_{t}$ is defined in Lemma C.2, a key term in the analysis becomes $P_{t}=-\mathbb{E}\|\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t-1};\xi_{t})\|_{F}^{2}\cdot A_{t}^{2}$, which is strictly negative. This leads to a tighter convergence bound, as it effectively introduces an additional beneficial term into the recurrence. However, computing this adaptive $\gamma_{t}$ is often impractical as it depends on quantities that are difficult to estimate during training. Consequently, we adhere to the common and more practical approach of using a constant $\gamma\leq 1$.
@@ 4.2 Non-Ergodic Convergence of Muon
The PL condition has been widely employed in the convergence analysis of various first-order algorithms Karimi et al. (2016); Xie et al. (2020a); Li & Li (2022), though typically under restricted settings. Note that if $f(\cdot)$ is strongly convex, then it is necessarily convex and satisfies the PL condition. However, the converse does not hold in general; a counterexample is given by $f(x)=x^{2}+3\sin(2x)$.
@@
Suppose Assumptions 4.1, 4.2, 4.4, and 4.5 hold. Let $\{\mathbf{X}_{t}\}$ be the sequence of iterates generated by Algorithm 1 with a step size of $\eta_{t}=t^{-3/4}$. For either of the following options:
@@
    For Option MVR1 ($\gamma=0$), using the update rule (4) with momentum $\beta_{t}=1-t^{-1/2}$,
@@
    For Option MVR1 ($\gamma_{t}=t^{-1/2}$), using the update rule (2) with parameters $\beta_{t}=1-(t+1)^{-1/2}$,
the expected suboptimality gap after $T$ iterations is bounded by:
$\mathbb{E}[f(\mathbf{X}_{T+1})]-f^{*}\leq\mathcal{O}\left(\frac{(\ln T)^{2}}{T^{1/2}}\right).$
@@
Suppose Assumptions 4.1, 4.2, 4.4, and 4.5 hold. Let $\{\mathbf{X}_{t}\}$ be the sequence of iterates generated by Algorithm 1 with Option MVR2 (see update rule (5)). By setting the parameters to $\eta_{t}=t^{-2/3}$, $\beta_{t}=1-t^{-2/3}$, and $\gamma=1$, the expected suboptimality gap after $T$ iterations is bounded by:
$\mathbb{E}[f(\mathbf{X}_{T+1})]-f^{*}\leq\mathcal{O}\left(\frac{(\ln T)^{2}}{T^{2/3}}\right).$
@@
The convergence bounds in Theorems 4.3, and 4.4 are asymptotic, holding for any number of iterations $T$ exceeding a sufficiently large threshold $T_{A}$. The existence of this threshold is established in our proofs and depends on the selection of a helper constant $A$. The specific conditions on $A$ and definitions of $\mathcal{A}$ are as follows:
  • For Theorem 4.3 (MVR1 with $\gamma=0$): $A>\frac{\mathcal{A}}{8C}$ with $\mathcal{A}=6L^{-1}\sigma^{2}+8\sqrt{2}Ln+2Ln+L$.
  • For Theorem 4.3 (MVR1): $A>\frac{\mathcal{A}}{8C}$ with $\mathcal{A}=14L^{-1}\sigma^{2}+16\sqrt{2}Ln+2Ln+L$.
  • For Theorem 4.4 (MVR2): $A>\frac{\mathcal{A}}{6C}$ with $\mathcal{A}=20L^{-1}\sigma^{2}+34Ln+L$.
In all cases, the constant $C$ is defined as $C=\frac{\sqrt{2\mu}}{2}$. The proofs of Theorems 4.3, and 4.4 are provided in Appendix F, and G, respectively.
Remark 4.6.
Theorem 4.4 achieves a faster convergence rate of $\tilde{\mathcal{O}}(T^{-2/3})$ than the $\tilde{\mathcal{O}}(T^{-1/2})$ in Theorem 4.3, owing to its more effective variance reduction. While our rates are not as fast as the $\mathcal{O}(1/T)$ rate for Adam He et al. (2023), our analysis is more general, as it does not require the bounded stochastic gradient assumption used therein. Closing this rate gap under our relaxed assumption is an important direction for future work.
Remark 4.7.
Our non-ergodic convergence proofs are unified by a key recursive inequality for a Lyapunov function $\Delta_{t}=\mathbb{E}[f(\mathbf{X}_{t})]-f^{*}$:
$\Delta_{t+1}\leq\Delta_{t}-\frac{\sqrt{2\mu}}{t^{p}}\sqrt{\Delta_{t}}+\Gamma_{t}.$
The final convergence rate is dictated by the exponent $p\in(0,1)$. We establish this with $p=3/4$ for Theorem 4.3. For Theorem 4.4, superior variance control allows proving the inequality with a tighter $p=2/3$, leading to the faster rate.
@@ 5 Experiments
The precise, Variance-Reduction formulation (5) is more rigorously aligned with stochastic optimization theory and, as demonstrated in our experiments, consistently achieves the highest validation accuracy. However, its computational overhead can be prohibitive in large-scale settings, as shown in Figure 1c. We note that prior analysis in Sfyraki & Wang (2025); Sato et al. (2025) suggests that increasing batch size can mitigate the impact of stochastic noise on convergence properties. Given that mini-batch training is the standard paradigm in practice, a trade-off between theoretical exactness and computational feasibility is often inherent. Therefore, echoing the sentiment in Yuan et al. (2024), we posit that the approximate version serves as a pragmatic alternative when computational efficiency is a primary concern, as it incurs only a minimal degradation in performance. Conversely, in scenarios where maximizing validation accuracy is paramount, the precise formulation remains the recommended approach.
@@ 6.1 The Evolution of Optimization Algorithms
One of the cornerstones of deep learning is first-order optimization methods. In early research, the Momentum method proposed by Nesterov (1983) accelerates the convergence process by accumulating historical gradient information. Subsequently, the RMSprop algorithm introduced by Hinton et al. (2012) pioneered the concept of adaptive learning rates, allowing different parameters to have independent update step sizes. The Adam optimizer, proposed by Kingma & Ba (2014), integrated these two ideas and has become one of the most widely used optimizers in the deep learning field by adaptively adjusting the first and second moments of the gradients.
Building upon Adam, the academic community has proposed a series of improvements to address its limitations in various scenarios. To enhance the stability of the optimization process, Reddi et al. (2018) proposed AMSGrad, which uses the maximum of historical second moments to constrain the learning rate update. Dozat (2016) integrated Nesterov momentum into Adam to design NAdam, aiming for superior performance. To address issues in Adam concerning weight decay and learning rate adjustment, Loshchilov (2017) developed AdamW, which improves the effectiveness of L2 regularization by decoupling weight decay. Luo et al. (2019) designed AdaBound, which sets dynamic bounds on the learning rate to avoid excessively large or small update steps. Furthermore, to mitigate variance fluctuations in the early stages of training, Liu et al. (2019) proposed RAdam to rectify the variance estimate. Zhuang et al. (2020) proposed AdaBelief, which adjusts the second moment by calculating the deviation between the gradient and its moving average, aiming to enhance the model’s generalization ability. In recent years, more novel Adam variants have been proposed. For instance, Xie et al. (2020b), from a dynamical systems perspective, constructed the Adai framework to accelerate training and improve the quality of convergence points by decoupling the effects of the adaptive learning rate and momentum. Xie et al. (2024) proposed Adan, which introduces a novel Nesterov momentum estimation method to accelerate convergence without incurring additional gradient computation overhead. The latest works, such as Liang et al. (2024), introduce a Cautious masking strategy to enhance optimization efficiency. Yuan et al. (2024) proposed MARS to further improve the variance-reduction strategy in Adam.
@@
Beyond direct improvements to Adam, researchers have also explored other efficient optimization paradigms. A significant branch among these is Shampoo-type methods based on preconditioning with parameter curvature. Shampoo, proposed by Gupta et al. (2018), is one of the pioneering works in this area. Inspired by its framework, Muon Jordan et al. (2024); Liu et al. (2025a) was proposed. This method adapts to parameter curvature by orthogonalizing gradient momentum via the Newton-Schulz iteration, thereby improving convergence dynamics. Recently, several adaptive variants of Muon have also emerged. AdaMuon Si et al. (2025) incorporates element-wise adaptivity but lacks theoretical convergence guarantees. For efficient training of large language models, COSMOS Liu et al. (2025b) integrates ideas from SOAP Vyas et al. (2025) and Muon, demonstrating stability and memory advantages in practice but similarly lacks a convergence proof. In the literature, other related preconditioning methods include ASGO An et al. (2025), PolarGrad Lau et al. (2025), and AdaGO Zhang et al. (2025), which introduces the Adagrad-Norm step size Ward et al. (2020) into a simplified version of Muon. Meanwhile, other high-performing optimizers not belonging to the Shampoo family also warrant attention, such as Sophia Liu et al. (2023), which improves second-moment estimation through efficient diagonal Hessian approximation and coordinate clipping, and Lion Chen et al. (2023), a lightweight optimizer that only tracks momentum and uses the sign function to normalize updates.

### v3

@@ 3 Convergence Analysis
We begin our analysis by situating it within the established context of first-order stochastic optimization. The $\mathcal{O}(T^{-1/4})$ iteration complexity is a well-known bottleneck for methods like SGD, and recent work by shen2025convergence confirmed this limitation for the standard Muon algorithm. Accordingly, our first step in this section is to analyze the Muon-MVR1 variant and formally reaffirm this baseline iteration complexity in Theorem 3.1. While the rate itself is standard, our primary contribution here lies in the novel and unified analytical framework we develop to prove it, as this framework will be instrumental for subsequent results.
@@
These assumptions are quite common Zhou2018OnTC; Chen2018OnTC; Huang2021SUPERADAMFA; Guo2021ANC; Li2023ConvergenceOA; Wang2023ClosingTG; 10586270; Yuan2024MARSUT.
@@ 3.1 Erogdic Convergence of Muon
In this subsection, we establish the ergodic convergence of Muon-MVR1 and Muon-MVR2.
@@ 3.1.1 Option MVR1
We begin our analysis with Option MVR1, a straightforward implementation of one-batch momentum-based variance reduction. The following theorem establishes its ergodic convergence rate, demonstrating that the algorithm converges to a stationary point at a rate of $\tilde{\mathcal{O}}(T^{-1/4})$ for specific choices of learning rate and momentum schedules.
@@
Suppose Assumptions 3.1, 3.2, and 3.4 hold. Consider Algorithm 1 with a learning rate of $\eta_{t}=t^{-3/4}$. The expected average squared norm of the gradient is bounded for the following options:
@@
Theorems 3.1 establishes that as $T\to\infty$, the leading terms diminish to zero, and the algorithm converges to a neighborhood of a stationary point. The size of this neighborhood is determined by the stochastic gradient variance, the learning rate, and the momentum parameter. Notably, this convergence bound is free of any non-vanishing additive error term that depends on the dimension $n$ Sato2025ConvergenceBA. We absorb the dimensional dependence into the $\mathcal{O}$-notation to define $\tilde{\mathcal{O}}$, which yields the following complexity for the algorithm:
$\min_{t=1,\dots,T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq\tilde{\mathcal{O}}(T^{-1/4}).$
@@ 3.1.2 Option MVR2
While prior work has established a iteration complexity of $\tilde{\mathcal{O}}(T^{-1/3})$ for variance-reduction in non-convex settings for various algorithms, such as SGD fang2018spider; Cutkosky2019MomentumBasedVR; zhou2020stochastic and Adam Huang2021SUPERADAMFA; Yuan2024MARSUT, the theoretical convergence properties of this technique when applied to the Muon optimizer have remained an open question. We bridge this theoretical gap in Theorem 3.2, which rigorously proves that Option MVR2 of Algorithm 1 (Muon-MVR2) achieves the same $\tilde{\mathcal{O}}(T^{-1/3})$ iteration complexity in the general non-convex setting.
@@
Under Assumptions 3.1, 3.3, and 3.4, consider Algorithm 1 with Option MVR2 and update rule (5) parameters set as $\beta_{t}=1-\eta_{t}$, $\eta_{t}=t^{-2/3}$ for $t\geq 1$, and $\gamma=1$. Then, the following bounds hold:
$\begin{aligned} \frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}&\leq\frac{\sqrt{4\sigma^{2}+(16L^{2}n+8\sigma^{2})(1+\ln T)}}{T^{1/3}}+\frac{G}{T^{1/3}},\end{aligned}$
@@
Theorem 3.2 establishes a significant theoretical advantage. In contrast to prior work sfyraki2025lions that requires a computationally expensive growing batch size ($b=\mathcal{O}(T^{1/3})$) to control gradient variance, our result shows that Muon-MVR2 achieves the same $\tilde{\mathcal{O}}(T^{-1/3})$ iteration complexity in a more practical, fixed mini-batch setting. This is made possible by the specific hyperparameter schedule, where the learning rate $\eta_{t}=t^{-2/3}$ and the momentum parameter $\beta_{t}=1-\eta_{t}$ are critically coupled. This schedule ensures a careful balance between optimization progress and noise suppression.
@@
As noted in Yuan2024MARSUT, a more sophisticated, adaptive setting for the variance-reduction parameter $\gamma$ can be employed. Specifically, by setting $\gamma=\gamma_{t}=1-\frac{A_{t}}{\beta_{t}}$, where $A_{t}$ is defined in Lemma C.2, a key term in the analysis becomes $P_{t}=-\mathbb{E}\|\nabla f(\mathbf{X}_{t};\xi_{t})-\nabla f(\mathbf{X}_{t-1};\xi_{t})\|_{F}^{2}\cdot A_{t}^{2}$, which is strictly negative. This leads to a tighter convergence bound, as it effectively introduces an additional beneficial term into the recurrence. However, computing this adaptive $\gamma_{t}$ is often impractical as it depends on quantities that are difficult to estimate during training. Consequently, we adhere to the common and more practical approach of using a constant $\gamma\leq 1$.
@@ 3.2 Non-Ergodic Convergence of Muon
The PL condition has been widely employed in the convergence analysis of various first-order algorithms karimi2016linear; xie2020linear; li2022simple, though typically under restricted settings. Note that if $f(\cdot)$ is strongly convex, then it is necessarily convex and satisfies the PL condition. However, the converse does not hold in general; a counterexample is given by $f(x)=x^{2}+3\sin(2x)$.
@@
Suppose Assumptions 3.1, 3.2, 3.4, and 3.5 hold. Let $\{\mathbf{X}_{t}\}$ be the sequence of iterates generated by Algorithm 1 with a step size of $\eta_{t}=t^{-3/4}$. We analyze the following two MVR1 schemes:
@@
    Scheme 1 ($\gamma=0$): Using the update rule (3) with momentum $\beta_{t}=1-t^{-1/2}$. For this case, we define the constant $\mathcal{A}_{1}$ as:
    $\mathcal{A}_{1}=6L^{-1}\sigma^{2}+(8\sqrt{2}+2)Ln+L.$
@@
    Scheme 2 ($\gamma_{t}=t^{-1/2}$): Using the update rule (2) with parameters $\beta_{t}=1-(t+1)^{-1/2}$. For this case, we define the constant $\mathcal{A}_{2}$ as:
    $\mathcal{A}_{2}=14L^{-1}\sigma^{2}+(16\sqrt{2}+2)Ln+L.$
Then, for either scheme $i\in\{1,2\}$, there exists a constant $T_{0}=2e^{4}$ such that for all iterations $T\geq T_{0}$, the expected suboptimality gap is bounded as follows:
$\mathbb{E}[f(\mathbf{X}_{T+1})]-f^{*}\leq\frac{\mathcal{A}_{i}^{2}}{8\mu}\cdot\frac{(\ln T)^{2}}{T^{1/2}}.$
@@
Suppose Assumptions 3.1, 3.2, 3.4, and 3.5 hold. Let $\{\mathbf{X}_{t}\}$ be the sequence of iterates generated by Algorithm 1 with Option MVR2 (see update rule (5)). By setting the parameters to $\eta_{t}=t^{-2/3}$, $\beta_{t}=1-t^{-2/3}$, and $\gamma=1$, we define the constant $\mathcal{A}_{3}$ as :
$\mathcal{A}_{3}=20L^{-1}\sigma^{2}+34Ln+L.$
Then, there exists a constant $T_{0}=2e^{3}$ such that for all iterations $T\geq T_{0}$, the expected suboptimality gap is bounded as follows:
$\mathbb{E}[f(\mathbf{X}_{T+1})]-f^{*}\leq\frac{2\mathcal{A}_{3}^{2}}{9\mu}\cdot\frac{(\ln T)^{2}}{T^{2/3}}.$
Detailed proofs for Theorems 3.3 and 3.4 are provided in Appendix F and G, respectively.
@@
Our non-ergodic convergence proofs are unified by a recursive inequality on the potential function $\Delta_{t}=\mathbb{E}[f(\mathbf{X}_{t})]-f^{*}$:
$\Delta_{t+1}\leq\Delta_{t}-\frac{\sqrt{2\mu}}{t^{p}}\sqrt{\Delta_{t}}+\Gamma_{t},$
where $\sum_{i=1}^{t}\Gamma_{t}\leq\mathcal{O}(\ln t)$. The exponent $p\in(0,1)$ dictates the final convergence rate. For Theorem 3.3, we establish this inequality with $p=3/4$, which leads to a rate of $\tilde{\mathcal{O}}(T^{-1/2})$. The superior variance control in Theorem 3.4 enables a tighter analysis with $p=2/3$, which in turn yields the accelerated convergence rate of $\tilde{\mathcal{O}}(T^{-2/3})$.
@@ 4 Experiments
The precise, Variance-Reduction formulation (5) is more rigorously aligned with stochastic optimization theory and, as demonstrated in our experiments, consistently achieves the highest validation accuracy. However, its computational overhead can be prohibitive in large-scale settings, as shown in Figure 1c. We note that prior analysis in sfyraki2025lions; Sato2025ConvergenceBA suggests that increasing batch size can mitigate the impact of stochastic noise on convergence properties. Given that mini-batch training is the standard paradigm in practice, a trade-off between theoretical exactness and computational feasibility is often inherent. Therefore, echoing the sentiment in Yuan2024MARSUT, we posit that the approximate version serves as a pragmatic alternative when computational efficiency is a primary concern, as it incurs only a minimal degradation in performance. Conversely, in scenarios where maximizing validation accuracy is paramount, the precise formulation remains the recommended approach.
@@ 5.1 The Evolution of Optimization Algorithms
The advancement of deep learning relies on first-order optimization methods. Nesterov1983AMF proposed the Momentum method, which leverages historical gradients to accelerate convergence. Subsequently, hinton2012lecture introduced RMSprop, enabling per-parameter adaptive learning rates. Kingma2014AdamAM then integrated these ideas in Adam, an optimizer that adapts using the first and second moments of the gradients, establishing it as a standard choice in the field.
Numerous variants have been proposed to address the limitations of Adam.  Reddi2018OnTC introduced AMSGrad to ensure a non-increasing learning rate, while  dozat2016incorporating created NAdam by incorporating Nesterov momentum. To improve regularization,  loshchilov2017decoupled developed AdamW, which decouples weight decay from the optimization step. Other methods focus on controlling the learning rate and variance:  Luo2019AdaptiveGM proposed AdaBound to dynamically clip learning rates;  Liu2019OnTV designed RAdam to rectify variance estimates in early training; and  Zhuang2020AdaBeliefOA created AdaBelief to adapt step sizes based on gradient belief. Recent developments continue this trend, such as  Xie2020AdaptiveID which decouples adaptation from momentum in the Adai framework, and  10586270 which introduced Adan with a novel Nesterov momentum estimation. The latest works from  Liang2024CautiousOI and  Yuan2024MARSUT further enhance efficiency with new masking and variance-reduction strategies, respectively. Additionally, the MGUP strategy proposed by chang2025mgup smooths zero masks into small non-zero values to alleviate potential non-convergence issues that the Cautious mask may cause in Adam.
@@
Beyond Adam variants, research has explored other paradigms such as preconditioning methods that use parameter curvature. Gupta2018ShampooPS pioneered this direction with the Shampoo optimizer. Building on this work,  jordan2024muon proposed Muon, which adapts to curvature by orthogonalizing gradient momentum. Subsequent variants emerged, such as AdaMuon si2025adamuon which adds element-wise adaptivity, and COSMOS liu2025cosmos which integrates ideas from SOAP vyas2025soap for large model training. While these methods showed practical benefits, they often lacked convergence proofs. To bridge this theoretical gap for LMO-based methods, Gluon riabinin2025gluon introduces a novel layer-wise smoothness assumption, providing convergence guarantees that align with the practical implementations of optimizers like Muon and Scion pethick2025training. In the literature, other related preconditioning methods include ASGO An2025ASGOAS, PolarGrad lau2025polargrad, and AdaGO Zhang2025AdaGradMM, which introduces the Adagrad-Norm step size ward2020adagrad into a simplified version of Muon. Meanwhile, other high-performing optimizers not belonging to the Shampoo family also warrant attention, such as Sophia Liu2023SophiaAS, which improves second-moment estimation through efficient diagonal Hessian approximation and coordinate clipping, and Lion Chen2023SymbolicDO, a lightweight optimizer that only tracks momentum and uses the sign function to normalize updates.